OpenStudy (anonymous):

Samantha has 180 marbles, some are plain and some are colored. If there are 32 more plain marbles than colored marbles, how many colored marbles does he have? Write the equation and show all work please.

OpenStudy (darkbluechocobo):

so 32+x=180

OpenStudy (darkbluechocobo):

Actually switch that x+32=180

OpenStudy (darkbluechocobo):

because 32 more than x

OpenStudy (anonymous):

could you put that all in one big answer with a bit more explanation

OpenStudy (darkbluechocobo):

xD ok

OpenStudy (anonymous):

thank you for all your work

OpenStudy (darkbluechocobo):

32+x=180 Because It is stating that there is 32 MORE marbles than plain which is like stating 32+x. and there is total 180 Marbles so final conclusions would be again 32+x=180

OpenStudy (darkbluechocobo):

o wait

OpenStudy (darkbluechocobo):

Do you have to solve it also?

OpenStudy (anonymous):

yes

OpenStudy (darkbluechocobo):

in that case

OpenStudy (darkbluechocobo):

x=180-32

OpenStudy (darkbluechocobo):

so x=148

OpenStudy (darkbluechocobo):

but the thing is

OpenStudy (darkbluechocobo):

180-148 to find the actual colored marbles

OpenStudy (darkbluechocobo):

So let me put that into one big paragraph

OpenStudy (darkbluechocobo):

To find the amount of colored marbles you must first subtract 32 from 180 which is 148. Then you take 180-148 which equals 32

OpenStudy (darkbluechocobo):

xD i just realized sorry

OpenStudy (darkbluechocobo):

32+148=180

OpenStudy (anonymous):

its all good

OpenStudy (darkbluechocobo):

So that means the answer checks out so there you go :3

OpenStudy (darkbluechocobo):

It has been a pleasure helping you out my dear friend